Q4:- Suppose for a commodity X, demand and supply curves are given by the equations

                             (1)         Qd = 4 - P

                             (2)         Qs = -2+P

(a)  What are the price elasticities of the demand curve and the supply curves at the points P = 1, 2, 3, etc.?

(b)  Suppose for commodity X, the price elasticity of demand is constant throughout the curve. What can you say about the functional form of the demand curve?

1
Expert's answer
2021-07-29T06:54:01-0400


a)

Price elasticity of demand.

1*(1/4)=0.25. When P=1

2*(1/4)=0.5. When P=2

3*(1/4)=0.75. When P=3


Interpretation; the demand is inelastic


Price elasticity of supply

1*(1/2)= 0.5. When P=1

Inelastic


2*(1/2)=1. When P=2

Unitary


3*(1/2) =1.5. When P= 3

Elastic.


b)

It will show a demand curve with constant unit elasticity.